# Quantum sampling on a quantum annealer for large volumes in the strong coupling limit for gauge group U(3)

### v2

@@ I Introduction
Quantum annealers are promising devices for ensemble generation in statistical physics and field theory. The examples of quantum sampling range from boson sampling [3] to sampling proteins [4]. Also, many different methods are on the market for quantum annealers: the noisy Gibbs sampler [5], methods to investigate phase transition [6], hybrid approach [7], and the annealer as a quantum thermal sampler [8]. Quantum sampling of statistical ensembles requires however an additional step to obtain ensembles according to the target equilibrium distribution that depends on the physical parameters of the investigated theory [9, 10, 11]. On a classical computer, via a Monte Carlo, importance sampling is realized e.g. by the Metropolis algorithm. In our effective theory, which is governed by a constraint on admissible configurations, ensemble generation is a difficult task, specifically at low temperature. In contrast, on a quantum annealer, the admissible configurations can be generated easily with the so-called QUBO formalism which is independent of the temperature.
The scope of this paper is to address larger volumes compared to our previous studies [1, 2]. We use the D-wave advantage system that has the Pegasus topology [12] and consists of 5600 physical qubits. This system can only accommodate small lattice volumes. Large volumes, on the other hand, require an alternative scheme that can sample smaller sub-volumes, or sub-lattices. The central idea of our work is to make use of an iterative scheme that embeds sub-lattices in parallel with specific boundaries that are fixed by the previous quantum sampling step. This scheme ultimately allows us to generate larger volume configurations.
@@ II Effective theory of Lattice QCD
Our effective theory is derived from staggered fermions in the strong coupling limit, which is well known to have a dual representation in terms of integer-valued dual variables [13]. These dual variables consist of so-called monomers $m(n)$ and dimers $d_{\nu}(n)$ (with $n$ a lattice coordinate) for ${\rm U}(3)$ gauge theory, and also baryonic world-lines, which have to be included for ${\rm SU}(3)$ gauge theory. We have studied ${\rm SU}(3)$ in [2] but in this work we restrict ourselves to ${\rm U}(3)$ for simplicity. As explained detail in [1], the integer-valued dual variables can be mapped to a binary vector $x$ which can be adapted for the quadratic unconstrained binary optimization (QUBO) formalism:
@@
$b$ $=(3-d_{ext}^{(1)},3-d_{ext}^{(2)},3-d_{ext}^{(3)},3-d_{ext}^{(4)})\,,$ (17)
where $d_{ext}^{(x)}$ corresponds to the number of external dimers. This quantity changes after each update as explained in Sec. III.1.
@@ III Technical development
In order to sample our effective theory on D-wave, we have to map logical qubit $x$ as in Eq. II to the physical qubits with connectivities given by the Pegasus topology. This can be achieved by synchronized combination of physical qubits (”chain”) to form a logical qubit, as discussed in detail in [1]. With 5600 physical qubits available, only small volumes can be sampled at once.
For larger volumes, we use an iterative scheme that only samples disconnected sub-lattices, keeping dimers that are not part of the sub-lattice fixed, see Eq. 17. An example of the decomposition to sub-lattices is given in Fig. 1. Only the degrees of freedom inside the red boxes are sampled, and this is done in parallel to maximize the use of qubits.
Figure 1: Selecting $2\times 2$ sub volume with fixed boundary condition. Black dots are monomers, blue lines are dimers. Monomers and dimers inside of red boxes are updated in parallel. The Grassmann constraint is required to be satisfied at each site.
@@ III.1 Parallel Sampling scheme
In order to sample all degrees of freedom, we scan all possible sub-lattices, defined by the base point and all directions to span the volume by the sub-lattices, as shown in Table 1.
@@
We choose the sub-lattices of size $2\times 2$ (square sampling), and of size $2\times 2\times 2$ (cube sampling). Clearly, the sub-lattice dimension $d_{s}$ has to be smaller or equal to the dimension $d$ of the lattice volume. For $d_{s}<d$, we have to loop through $\binom{d}{d_{s}}$ sets of directions to sample the full volume. For $2\times 2$ sub-lattices, 16 logical qubits are required per sub-lattice and for $2\times 2\times 2$ sub-lattice, 40 logical qubits. Since the number of sub-lattices is $V/V_{sub}$, the number of logical qubits for lattice size $V=N_{s}^{(d-1)}\times N_{t}$ is $2V(1+d/2)$ with $d$ the dimension of the volume. If the lattice is too large, we choose randomly a maximal subset of sub-lattices ($2\times 2$: 170, $2\times 2\times 2$: 45) to sample in parallel. This results in a block-diagonal form of the QUBO matrix $Q$. The quantum annealer then solves for $x$ in a fixed amount of time regardless of how many qubits are used in the parallelization.
@@
The quantum annealer produces a distribution of valid configurations that depend on the penalty factor $p$ of the QUBO matrix defined in Eq. (II), as well as hyper parameters that we discuss in detail in the next section. We classify the valid solution vectors per sub-lattice into histograms for each set of boundary conditions $b$, and sort them by monomer number $M$ and number of temporal dimers $D_{t}$, which encode the quark mass and temperature dependence. As the set of generated histograms $h_{p}(b,M,Dt)$ will depend on both the physical parameters and in particular the penalty factor $p$, the histogram is an estimate of the true distribution $h_{true}(b,M,Dt)$ that we obtain from exact enumeration (see Fig. 2 and App. A. We will discuss in Sec. III.3 and III.4 two methods to correct the histograms to obtain valid results.
@@
Figure 4: Comparison for $2\times 2$ sub-lattices between exact histograms from enumeration with those generated by D-wave for $p=1$ and $p=2$. Left: Comparison of the multiplicities for 256 different boundaries. Right: Comparison of the weights for a specific boundary (3,3,3,3) for $am_{q}=1.0$ and $\gamma=0.1$, with exact Boltzmann distribution compared to approximate distributions from D-Wave.
In Fig. 4, a comparison of the true histogram $h_{true}(b)$ by exact enumeration and the histogram $h_{p}^{b}$ obtained by the D-Wave quantum annealer for penalty factors $p=1,2,\infty$ is shown. Here, $b$ labels the 256 distinct boundary conditions $b=\{b_{1},\dots b_{4}\}$ in lexicographical order (left). The distribution for a specific boundary condition $b=(3,3,3,3)$ (right) is shown for quark mass $am_{q}=1.0$ and $\gamma=0.1$, now sorted by the Boltzmann weight $e^{-S}$. Both histograms are in good agreement for $p=1$, but this agreement deteriorates for larger $p$. This indicates that we have good overlap between the true and the approximate distributions of the $2\times 2$ sub-lattices, which is important to implement our strategy for large lattice volumes. How to correct the distribution for the small deviations is discussed in Sec. III.3 and III.4.
@@ III.2 Optimizing hyper parameters
To maximize the efficiency of the simulation and obtain correct results, the hyper parameters chain_strength, num_reads, annealing_time should be optimized. Since the mapping from logical qubit to physical qubit is not trivial (we use automatized embedding which uses on average less than 2 physical qubits per logical qubit) it can happen that during the annealing process the embedding becomes invalid and results in broken ”chains”. For submitting a problem to the D-Wave system, one has to provide a value for the chain_strength, which can be chosen to reduce the number of broken chains. However, we cannot arbitrarily increase the value of chain_strength such that no broken chains occur, since too large a value adversely affects the number of obtained valid solutions. Thus some optimization is required.
@@
The QUBO matrix is rescaled by $Q_{max}$, which is the maximum absolute value of any element of the QUBO matrix before submitting to the D-wave solver. This results in rescaling the chain_strength to order one. For small values of chain_strength, the number of broken chains is large. As shown in Fig. 5, this results in a small validity rate. With increasing chain_strength, the validity rate grows with maximum around chain_strength is 1.0 to 1.2, depending on penalty factor $p$. Note that the weight matrix $W$ in the QUBO matrix has a block-diagonal structure, whereas the constraint matrix $A$ has long-range connectivity, requiring a larger chain_strength. However, if it is much larger than 1, the validity rate drops again [1].
@@
Figure 5: The validity rate $v$ as a function of chain_strength on $2\times 2$ sub-lattice with ’3 3 3 3’ boundary condition. For penalty factor $p=1,2$ and infinite. $\gamma=0.1,1$, $am_{q}=0.1$. Right: $\gamma=0.1$, $am_{q}=0.1$
@@ III.3 Metropolis and Metropolis-Hastings
Although we know the exact distributions of the $2\times 2$ sub-lattices from enumeration, this is of no help with extending to finite periodic lattices: First, analytically extending to larger sub-lattices such as $4\times 4$ is not feasible due the exponential growth of computation involved. Second, a standard Metropolis algorithm attempting to glue the sub-lattices together would produce extremely small acceptance rates of about $0.1\%$ independent of the physical parameters. Alternative classical algorithms such as the Metropolis algorithm using the parallelization scheme as in Fig. 1 and sweeping through the 4 base points gives acceptance rates of about 20% for low temperatures $\gamma=0.1$, with a quark mass dependence as show in Fig. 6 (right).
@@
Figure 6: Comparison of the acceptance rate. Left: for a $4\times 4$ lattice from Metropolis-Hastings, with $2\times 2$ sub-lattice histograms from D-Wave, for various penalty factors $p$. Center: illustrating independence of the acceptance rate of the volume. Right: Metropolis from classical computation, for $\gamma=0.1,1.0$. Metropolis-Hastings for $p=\infty$ reproduces the classical Metropolis acceptance rate.
In contrast, with the approximate distribution for $2\times 2$ sub-lattices measured by D-Wave for penalty factor $p=1$, we cannot incorporate the standard Metropolis algorithm, but have to use the Metropolis-Hastings algorithm:
@@
with $e^{-S_{\rm new}}$ and $e^{-S_{\rm old}}$ the Boltzmann factors for the new/old configuration, and $h_{\rm new}$ and $h_{\rm old}$ the new/old histograms of the configuration. They are non-trivial: although they have the same boundary condition $b$, they differ in the specific distribution of monomers and dimers within the $2\times 2$ sub-lattice. The proposal probability $P_{\rm proposal}=\frac{h_{\rm old}}{h_{\rm new}}$ for a set of $2\times 2$ updates is not uniform as it is drawn from the measured histogram. As pointed out in Sec. III.1, these histograms have good overlap with the true distribution. Hence we expect that the Metropolis-Hastings algorithm will have a much larger acceptance rate, in particular also for low temperatures. The dependence of the acceptance probability on the penalty factors $p$ is shown in Fig. 6 (left). We can clearly see that $p=1$ has the largest acceptance rate, and that large $p$ will decrease the acceptance rate. In the limit $p\rightarrow\infty$, the weight matrix will not contribute to the QUBO matrix and the $2\times 2$ histograms become flat. Hence, due to $h_{old}=h_{new}$, the proposal probability drops out in the Metropolis-Hastings, which simplifies to standard Metropolis. The acceptance rate will hence be minimal as there is poor overlap with the important configurations given by the physical parameters. As shown in Fig. 6 (c), the $p=\infty$ acceptance rate agrees with the classical Metropolis algorithm. Recall that classical computations are in particular expensive for low temperatures (small $\gamma$), whereas the computational costs for the histograms measured by D-wave do not depend on $\gamma$. Also, the acceptance rate does not depend much on the volume as shown in Fig. 6 (b), as expected.

### v3

@@ I Introduction
Quantum annealers are promising devices for ensemble generation in statistical physics and field theory. The examples of quantum sampling range from boson sampling [3] to sampling proteins [4]. Also, many different methods are on the market for quantum annealers: the noisy Gibbs sampler [5], methods to investigate phase transition [6], hybrid approach [7], and the annealer as a quantum thermal sampler [8]. Quantum sampling of statistical ensembles requires however an additional step to obtain ensembles according to the target equilibrium distribution that depends on the physical parameters of the investigated theory [9, 10, 11]. In contrast to gate-based quantum computing, quantum annealers are restricted to optimization problems, and a crucial step is to map the physical system on the qubits of the annealer. This embedding poses a scalability challenge [12, 13], which limits the usefulness for realistic applications further. Despite these challenges, some progress in high energy physics such as scalar field theory [14], lattice gauge theory with non-Abelian gauge groups ${\rm SU}(2)$ [15], dihedral group $D_{3}$ [16], and for building blocks of standard model physics  [17] has been made. We address in this paper also a non-Abelian gauge group ${\rm U}(3)$, which at strong coupling has similar properties as ${\rm SU}(3)$ at vanishing quark density, as explained in Sec. II.
On a classical computer, via a Monte Carlo, importance sampling is realized e.g. by the Metropolis algorithm. The effective theory which we study is based on the dual formulation of lattice QCD at strong coupling, and is governed by a constraint on the dual variables, which restricts admissible configurations. Depending on the physical parameters, ensemble generation is a difficult task, specifically at low temperature where the entropy of the system grows rapidly. In contrast, on a quantum annealer, the admissible configurations can be generated easily with the so-called QUBO formalism which scales independently of the physical parameters such as the temperature, as explained in Sec. II.
The scope of this paper is to address larger volumes compared to our previous studies [1, 2], which requires new strategies to take advantage of the quantum annealer. We use the D-Wave advantage system with the Pegasus topology [18] that has 5760 physical qubits and is the most advanced annealer up to date. In Fig. 2 (left) we show its connectivity. This system can however only accommodate small lattice volumes. Large volumes, on the other hand, require an alternative scheme that samples smaller sub-lattices. The central idea of our work is to make use of an iterative scheme that embeds sub-lattices in parallel with specific boundaries that are fixed by the previous quantum sampling step. This scheme ultimately allows us to generate larger volume configurations.
@@ II Effective theory of Lattice QCD
Our effective theory is derived from staggered fermions in the strong coupling limit, which is well known to have a dual representation in terms of integer-valued dual variables:
$S_{F}$ $=\sum_{x,\nu}\frac{\gamma^{\delta_{\nu 0}}}{2}\eta_{\nu}(x)\big{(}e^{\mu\delta_{\nu 0}}\overline{\chi}(x)U_{\nu}(x)\chi(x+\hat{\nu})-e^{-\mu\delta_{\nu 0}}\overline{\chi}(x+\hat{\nu})U^{\dagger}_{\nu}(x)\chi(x)\big{)}+am_{q}\sum_{x}{\bar{\chi}_{x}\chi_{x}}$ (1)
$Z$ =$=\sum_{\{d,m\}}\prod_{b=(n,\hat{\nu})}\frac{(N_{c}-d_{\nu}(n))!}{N_{c}!d_{\nu}(n)!}\gamma^{2d_{\nu}(n)\delta_{\hat{0},\hat{\nu}}}\prod_{n}\frac{N_{c}!}{m(n)!}(2am_{q})^{m(n)}\,,$ (2)
where Eq. (2) is obtained from Eq. (1) by integrating out the gauge variables first, and afterwards the Grassmann variables [19]. As we restrict to ${\rm U}(3)$ gauge group here, baryons are absent in Eq. (2), they would form self-avoiding world-lines for ${\rm SU}(3)$. We have studied ${\rm SU}(3)$ in [2] but in this work we restrict ourselves to ${\rm U}(3)$ for simplicity. The dual variables in Eq. (2) consist of so-called monomers $m(n)$ and dimers $d_{\nu}(n)$ (with $n$ a lattice coordinate) Dimers correspond to pion hops, for ${N_{c}}=3$ there can be up to 3 pions on the same link (due to the quark sub-structure), the monomers contribute to the chiral condensate. On each site, the number of monomers and dimers touching this site has to add up to ${N_{c}}$ due to the Grassmann constraint.
This system describes pions that are strongly interacting at low temperatures, resulting in a large entropy, and non-interacting at high temperatures. The chiral transition is located where the pion correlation is maximal (and diverges in the chiral limit). The temperature (in lattice units) is given by $aT=\frac{\gamma^{2}}{N_{t}}$ where $\gamma$ is an anisotropy that favors temporal pion hops over spatial hops for $\gamma>1$, and vice versa for $\gamma<1$.
As explained detail in [1], the integer-valued dual variables can be mapped to a binary vector $x$ which can be adapted for the quadratic unconstrained binary optimization (QUBO) formalism:
@@
$b$ $=(3-d_{ext}^{(1)},3-d_{ext}^{(2)},3-d_{ext}^{(3)},3-d_{ext}^{(4)})\,.$ (19)
The boundary $b$ is no longer constant as in our previous work, but now depends on the number of external dimers $d_{ext}^{(x)}\in\{0,1,2,3\}$. This quantity changes after each update as explained in Sec. III.1.
@@ III Technical development
In order to sample our effective theory on D-Wave, we have to map logical qubit $x$ as in Eq. (II) to the physical qubits with connectivities given by the Pegasus topology. This can be achieved by synchronized combination of physical qubits (”chain”) to form a logical qubit, as discussed in detail in [1]. With 5760 physical qubits available, only small volumes can be sampled at once.
(a) Pegasus Topology
(b) Embedding of 4 sub-lattices
Figure 1: Left: the connectivity between qubits of the Pegasus topology, showing a part of the full 5760 qubits. Qubits in this topology are considered to have a nominal length of 12 (each qubit is connected to 12 orthogonal qubits through internal couplers) and degree of 15 (each qubit is coupled to 15 different qubits) [^2] https://docs.dwavesys.com/docs/latest/c_gs_4.html. Right: embedding of sub-lattices in parallel, here shown for a subset of 4, using automatic embedding, illustrating that each sub-lattice is disconnected. Each sub-lattice uses 24 physical qubits. Each monomer and dimer needs 2 logical qubits to represent integers between 0 and 3.
For larger volumes, we use an iterative scheme that only samples disconnected sub-lattices, keeping dimers that are not part of the sub-lattice fixed, see Eq. (19). An example of the decomposition to sub-lattices is given in Fig. 2. Only the degrees of freedom inside the red boxes are sampled, and this is done in parallel to maximize the use of qubits.
Figure 2: Selecting $2\times 2$ sub volume with fixed boundary condition. Black dots are monomers, blue lines are dimers. Monomers and dimers inside of red boxes are updated in parallel. The Grassmann constraint is required to be satisfied at each site. The dimers $d_{ext}$ outside of red boxes are external dimers (light blue) and give boundary condition for sampling the red sub-lattices.
@@ III.1 Parallel Sampling scheme
We now explain the details of the parallel sampling scheme, leaving the details on the Monte Carlo updates for Sec. III.3. In order to sample all degrees of freedom, we have to alternate which dimers are frozen and which are updated. This we do by iterating through the base-points, as high-lighted in Fig. 2, which define what array of sub-lattices have to be considered for the next update. The number of base-points depend on the dimension and the sub-lattice structure, as shown in Table 1.
@@
We choose the sub-lattices of size $2\times 2$ (square sampling), and of size $2\times 2\times 2$ (cube sampling). Clearly, the sub-lattice dimension $d_{s}$ has to be smaller or equal to the dimension $d$ of the lattice volume. For $d_{s}<d$, we have to loop through $\binom{d}{d_{s}}$ sets of directions to sample the full volume. For $2\times 2$ sub-lattices, 16 logical qubits($\sim 24$ physical qubits) are required per sub-lattice and for $2\times 2\times 2$ sub-lattice, 40 logical qubits($\sim 88$ physical qubits). Since the number of sub-lattices is $V/V_{sub}$, the number of logical qubits for lattice size $V=N_{s}^{(d-1)}\times N_{t}$ is $N_{\rm logical}=2V(1+d/2)$, with $d$ the dimension of the volume and $N_{s}$ and $N_{t}$ are spatial and temporal size of lattice, respectively. Due to the block-diagonal structure of the QUBO matrix $Q$, the embedding of $N_{\rm logical}$ logical qubits to the physical qubits results in disconnected sub-lattice embeddings as shown in Fig. 2 (right). The quantum annealer then solves for $x$ in a fixed amount of time regardless of how many qubits are used in the parallelization. Hence, as long as the lattice volume $V$ fits on the annealer at once (for $2\times 2$: 170 sub-lattices, $2\times 2\times 2$: 45 sub-lattices) the scaling is constant. If the lattice is too large, we sequentially update a block of maximal subset of sub-lattice at a time. In that case, sampling the sub-lattices scales linear with the volume.
@@
The quantum annealer produces a distribution of valid configurations that depend on the penalty factor $p$ of the QUBO matrix defined in Eq. (II), as well as hyper parameters that we discuss in detail in the next section. We classify the valid solution vectors per sub-lattice into histograms for each set of boundary conditions $b$, and sort them by monomer number $M$ and number of temporal dimers $D_{t}$, which encode the quark mass and temperature dependence. As the set of generated histograms $h_{p}(b,M,Dt)$ will depend on both the physical parameters and in particular the penalty factor $p$, the histogram is an estimate of the true distribution $h_{true}(b,M,Dt)$ that we obtain from exact enumeration (see Fig. 3 and App. A. We will discuss in Sec. III.3 and Sec. III.4 two methods to correct the histograms to obtain valid results.
@@
Figure 5: Comparison for $2\times 2$ sub-lattices between exact histograms from enumeration with those generated by D-Wave for $p=1$ and $p=2$. Left: Comparison of the multiplicities for 256 different boundaries. Right: Comparison of the weights for a specific boundary $b=$(3,3,3,3) for $am_{q}=1.0$ and $\gamma=0.1$, with exact Boltzmann distribution $h_{true}(b)$ compared to approximate distributions from D-Wave $h_{p}(b)$.
In Fig. 5, a comparison of the true histogram
$h_{true}(b)$ $=\frac{e^{-S[C_{2\times 2}(b)]}}{\sum\limits_{C_{2\times 2}(b)}e^{-S[C_{2\times 2}(b)]}}$ (21)
by exact enumeration (with $C_{2\times 2}(b)$ the $2\times 2$ configurations at fixed boundary $b)$ and the histogram $h_{p}^{b}$ obtained by the D-Wave quantum annealer for penalty factors $p=1,2,\infty$ is shown. Here, $b$ labels the 256 distinct boundary conditions $b=\{b_{1},\dots b_{4}\}$ in lexicographical order (left). The distribution for a specific boundary condition $b=(3,3,3,3)$ (right) is shown for quark mass $am_{q}=1.0$ and $\gamma=0.1$, now sorted by the Boltzmann weight $e^{-S}$. Both histograms are in good agreement for $p=1$, but this agreement deteriorates for larger $p$. This indicates that we have good overlap between the true and the approximate distributions of the $2\times 2$ sub-lattices, which is important to implement our strategy for large lattice volumes. How to correct the distribution for the small deviations is discussed in Sec. III.3 and Sec. III.4.
@@ III.2 Optimizing hyper parameters
To maximize the efficiency of the simulation and obtain correct results, the hyper parameters chain_strength, num_reads, annealing_time should be optimized. As described earlier, we must embedded logical qubits into the physical layout of qubits so that the constellation of our logical qubits can simulate our QUBO problem (we use D-Wave’s automatized embedding to do this mapping). The physical qubits comprising a logical qubit must act as one; either all take values of 0 or 1. And their connections with other logical qubits must also act accordingly. However, during the annealing process it may (and does) happen that physical qubits within a logical qubit do not act uniformly, and this results in broken “chains”. These broken chains result in invalid solutions.
For submitting a problem to the D-Wave system, one has to provide a value for the chain_strength, which acts as a constraint that enforces connectivity of the chains. Increasing this value can reduce chain breaks. However, we cannot arbitrarily increase the value of chain_strength such that no broken chains occur, since too large a value makes our QUBO matrix less relevant and the system approaches the state of $n$ independent logical qubits that do not interact with each other. This in turn adversely affects the number of obtained valid solutions. Thus some optimization is required.
@@
We rescale the QUBO matrix manually by $Q_{max}$, which is the maximum absolute value of any element of the QUBO matrix, before submitting to the D-Wave solver. This results in the rescaling of the chain_strength to order one. For small values of chain_strength, the number of broken chains is large. Since broken chains produce almost always invalid configurations, as a result the validity rate is low, as shown in Fig. 6. With increasing chain_strength, the validity rate grows with maximum around chain_strength is between 1.0 to 1.4, depending on the penalty factor $p$ and physical parameters. Note that the weight matrix $W$ in the QUBO matrix has a block-diagonal structure, whereas the constraint matrix $A$ has long-range connectivity, requiring a larger chain_strength. For larger chain_strength the validity rate drops again.
@@
Figure 6: Left: The validity rate $v$ as a function of chain_strength on $2\times 2$ sub-lattice with ’3 3 3 3’ boundary condition. For penalty factor $p=1,2$ and infinite. $\gamma=0.1,1.0$, $am_{q}=0.1$. Right: The validity rate as function of the annealing time for $p=1,2$ and infinite and $\gamma=0.1$, $am_{q}=0.1$. A plateau is reached after about 100ms.
@@ III.3 Metropolis and Metropolis-Hastings
Although we know the exact distributions of the $2\times 2$ sub-lattices from enumeration in this simplest case (see App. A), this is not the case in the more realistic cases in general (SU(3) including strong coupling expansion). But even with this exact distribution, it is not feasible to extend to finite periodic lattices: First, analytically extending to larger sub-lattices such as $4\times 4$ can not be achieved due the exponential growth of computation involved. Second, a standard Metropolis algorithm attempting to glue the sub-lattices together would produce extremely small acceptance rates of about $0.1\%$ independent of the physical parameters. Alternative classical algorithms such as the Metropolis algorithm using the parallelization scheme as in Fig. 1 and sweeping through the 4 base points gives acceptance rates of about 20% for low temperatures $\gamma=0.1$, with a quark mass dependence as shown in Fig. 7 (right).
@@
Figure 7: Comparison of the acceptance rate versus $am_{q}$. Left: for a $4\times 4$ lattice from Metropolis-Hastings, with $2\times 2$ sub-lattice histograms from D-Wave, for various penalty factors $p$. Center: illustrating independence of the acceptance rate of the volume. Right: Metropolis from classical computation, for $\gamma=0.1,1.0$. Metropolis-Hastings for $p=\infty$ reproduces the classical Metropolis acceptance rate.
In contrast, with the approximate distribution for $2\times 2$ sub-lattices measured by D-Wave for penalty factor $p=1$, we do not rely at all on the exact distribution obtained from classical computation at all. However, we cannot incorporate the standard Metropolis algorithm, but have to use the Metropolis-Hastings algorithm:
@@
with $e^{-S_{\rm new}}$ and $e^{-S_{\rm old}}$ the Boltzmann factors for the new/old configuration, and $h_{\rm new}$ and $h_{\rm old}$ the new/old histograms of the configuration. They are non-trivial: although they have the same boundary condition $b$, they differ in the specific distribution of monomers and dimers within the $2\times 2$ sub-lattice. The proposal probability $P_{\rm proposal}=\frac{h_{\rm old}}{h_{\rm new}}$ for a set of $2\times 2$ updates is not uniform as it is drawn from the measured histogram. The details of the procedure are described by the following steps:
  1. 1.
    On the classical computer, generate a large lattice filled with 3 monomers at every sites.
  2. 2.
    Select independent sub-lattices and use the set of boundaries $\{b\}$ to construct the QUBO matrix for the physical parameters.
  3. 3.
    Generate a distribution for the sub-lattices using the quantum annealer and store it on the memory of classical computer. If the distribution for a specific boundary condition already exists, we can skip this step.
  4. 4.
    Select for each sub-lattice $h_{new}$ from the distribution and make an accept/reject step by comparing with the previous configuration $h_{old}$ on the classical computer according Eq. (23).
  5. 5.
    Store the configuration in terms of monomers and dimers and measure observables.
  6. 6.
    Change the base point. Repeat from step 2.
As pointed out in Sec. III.1, these histograms have good overlap with the true distribution. Hence we expect that the Metropolis-Hastings algorithm will have a much larger acceptance rate, in particular also for low temperatures. The dependence of the acceptance probability on the penalty factors $p$ is shown in Fig. 7 (left). We can clearly see that $p=1$ has the largest acceptance rate, and that large $p$ will decrease the acceptance rate. In the limit $p\rightarrow\infty$, the weight matrix will not contribute to the QUBO matrix and the $2\times 2$ histograms become flat. Hence, due to $h_{old}=h_{new}$, the proposal probability drops out in the Metropolis-Hastings, which simplifies to standard Metropolis. The acceptance rate will hence be minimal as there is poor overlap with the important configurations given by the physical parameters. As shown in Fig. 7 (c), the $p=\infty$ acceptance rate agrees with the classical Metropolis algorithm. Recall that classical computations are in particular expensive for low temperatures (small $\gamma$), whereas the computational costs for the histograms measured by D-Wave do not depend on $\gamma$. This is particularly evident by comparing the statistical uncertainties of the chiral condensate as measured via the worm algorithm versus the Metropolis-Hastings results based on the D-Wave histograms, as shown in Fig. 8. In the case of worm simulation, the statistical uncertainty grows with decreasing temperature, while it remains constant for the results based on the hybrid approach of D-Wave histograms and Metropolis-Hastings. Interestingly, the error for $8^{3}\times 4$ is of comparable size as for $4^{3}\times 4$, both having the same number of Metropolis-Hastings updates. This might be related to the fact that for each base point, the solution vector $x$ scales with the volume, as it contains $2\times 2$ sub-lattices proportional to the volume.
Also, the acceptance rate does not depend much on the volume as shown in Fig. 7 (b), as expected.
Figure 8: The chiral condensate as a function of $\gamma$ which is related to the temperature $aT=\frac{\gamma^{2}}{N_{T}}$. While the size of error of classical simulations via worm algorithm depend on the physical parameter, the method that we demonstrate in this paper is independent of $\gamma$, as illustrated in the inset. Hence, we confirm that this method has a significant advantage over the classical method at low temperatures.